# AppSupplemental Materials References

### v1

@@ 8 Conclusion
Following the same staggered DiD design, we report the ATT estimates using the IRT estimates instead in Figure S.5.7. Similarly, the null effect after a shooting ($t=0$ and beyond) holds for all sets of comparisons.
@@
Figure S.5.7: ATT Estimates for Staggered DiD Demonstrate Null Results Across Party When Using Bayesian IRT Model
@@
S.6 Treatment Timing in California
In California, Legislators Bloom, Cortese, Feuer, Hurtado, Liu, and Min (the semi-transparent green observations) are dropped due to the fact that we do not have at least one pre-treatment and one post-treatment observation. Two other treated legislators, Assemblyman Mike Feuer and State Senator Carol Liu, are dropped because they were previously treated by the North Hollywood shooting (so we do not have a pre-treatment observation for them, only post-treatment). This can be seen in Figure S.6.1.
Figure S.6.1: Treatment Timing for Legislators Representing Districts in Which a Mass Shooting Occurred
@@
S.7 Treatment at Varying Distances
@@
To determine shootings that are affected under the varying distance, we calculate a 5 and 25 mile radii, in euclidean distance, around each shooting’s location. We use the North American Datum of 1983 (NAD83) geodetic datum as our Coordinate Reference System (CRS). Figure S.7.1 shows these buffers, in blue and green, respectively. Figure S.7.1 zooms in on Southern California and shows shootings occurring between 2012 and 2022 only.
Figure S.7.1: Treatment Assignment Strategy for Varying Distances
@@
S.8 Assigning Treatment to a District
To estimate the effects of a mass shooting on the Gun Control Score of a district, we first confine our period of study to 2012-2022 to remove any concerns that we are not appropriately accounting for redistricting. Our estimation strategy is similar to those underlying our analyses of individual legislators, though we include an interacted district- and chamber-fixed effect and cluster standard errors by district and chamber. The full results can be seen in Table S.8.1.
Table S.8.1: ATT Estimates for Effect of Mass Shooting on District
@@
# of District-Chambers 117 89 44
# of Mass Shootings 11 11 10
District x Chamber FE
Pre-Treatment Mean 0.056 0.343 -0.176
  • Note: Robust standard errors are clustered at the district and chamber level. Gun Control Score is measured from -1 to 1, from least to most supportive of restrictive firearms legislation. Table reports the estimated effects of mass shootings occurring within a district on the Gun Control Score of the district. Individual legislators who switch parties throughout the duration of study are removed.
S.9 Additional States
Figure S.9.1: Gun Control Score of State Legislators across 6 States by Party, 2011-2022
Note: Each black point is an individual legislator’s score for the corresponding year of the legislative session. Large colored points are binned averages of each party’s score, across states, in a given year. Linear regression smoothers (with dotted CIs) are fit to legislators in either party (and across parties, in black) to show general trends in partisan voting behaviors across session-years. The figure includes legislators in the upper and lower chambers of California (2011-2022), Colorado (2012-2022), Florida (2011-2021), Texas (biennial, 2013-2021), Missouri (2019-2022), and Georgia (biennial, 2011-2021). The “bridge” agents in these states are: Donald Coram (CO); David Ralston (GA); Debbie Mayfield (FL); Mike Bernskoetter (MO); Dan Huberty (TX).
S.10 Differential Effects of Shootings
Figure S.10.1: ATT Estimates for Effect of Mass Shooting on Gun Control Score by Post-Treatment Group Suggest Nuanced Results
Note: Figure plots the ATT and 95% Confidence Intervals for each treatment cohort in the first post-treatment period, $t=0$. Treatment cohorts are defined by the first year in which observations are considered post-treatment, and the y-axis reports the group while the x-axis plots the ATT.
As Figure S.10.1 shows, we find statistically significant positive effects for mass shootings on the Gun Control Scores of treated Republicans, relative to other Republicans, in groups 2019 and 2018, and a negative effect in 2016. We find statistically significant positive effects on treated Democrats in groups 2019 and 2017. We also find statistically significant positive effects on all legislators in 2018 and 2017, and a negative effect in 2016. Across all other groups, the estimates are not statistically significant, and many are quite noisy, further suggesting that there is variation in the effects of particular shootings within post-treatment period groups.
Our most precise estimate in Figure S.10.1 is for Democrats in group 2017 (ATT = 0.530, SE = 0.031, which suggests an 26.5% increase in votes for restrictive policies among treated Democrats), though this estimate does not represent the ATT of a particular shooting. Instead, it is the aggregated ATT for three distinct shootings: (1) a July 2016 shooting in Dallas, TX with 5 victims killed, (2) a January 2017 shooting in Fort Lauderdale, FL with 5 victims killed, and (3) a June 2017 shooting in Orlando, FL with 5 victims killed.
[…]
S.11 Cosponsorship Network Analysis and Gun Policy Activity
S.11.1 Cosponsorship Network
Figure S.11.1: Cosponsorship Network Among California Legislators on Firearm-related Legislation
Table S.11.1: Legislators with Highest Centrality in Each Chamber and Session
@@
S.11.2 Alternative Behavior
Table S.11.2: ATT Estimates for Effect of Mass Shooting on Legislators’ Activity in Gun Policy Domain
@@
Leg. x Chamber FE
Pre-Treatment Mean 0.014 0.017 0.0.12 0.600 0.792 0.423
  • Note: Standard errors are clustered at the legislator level in parentheses. Table reports the estimated effects of mass shootings occurring within a legislator’s district on a legislator’s activity in the gun policy domain for periods $t=-1$ to 2. Activity is a measure of the number of bills sponsored or cosponsored by a legislator related to gun policy out of all bills related to gun policy in a given year. Outdegree is a measure of legislator’s level of activity in cosponsoring firearms-related legislation. Individual legislators who switch parties throughout the duration of study are also removed.

### v2

@@ A.6.1 Results Using IRT Estimates
Following the same staggered DiD design, we report the ATT estimates using the IRT estimates instead in Figure A.6.1. Similarly, the null effect after a shooting ($t=0$ and beyond) holds for all sets of comparisons.
@@
Figure A.6.1: ATT Estimates for Staggered DiD Demonstrate Null Results Across Party When Using Bayesian IRT Model
@@
A.6.2 Results Using Rankings
Following the same staggered DiD design, we report the ATT estimates using the relative rankings of legislators in California — grouping legislators by year and ranking their Gun Control Scores relative to one another within each year — in Figure A.6.2. Similarly, the null effect after a shooting ($t=0$ and beyond) holds for all sets of comparisons.
Figure A.6.2: ATT Estimates for Staggered DiD Demonstrate Null Results Across Party When Using Bayesian IRT Model
A.7 Treatment Timing in California
In California, Legislators Bloom, Cortese, Feuer, Hurtado, Liu, and Min (the semi-transparent green observations) are dropped due to the fact that we do not have at least one pre-treatment and one post-treatment observation. Two other treated legislators, Assemblyman Mike Feuer and State Senator Carol Liu, are dropped because they were previously treated by the North Hollywood shooting (so we do not have a pre-treatment observation for them, only post-treatment). This can be seen in Figure A.7.1.
Figure A.7.1: Treatment Timing for Legislators Representing Districts in Which a Mass Shooting Occurred
@@ A.7 Treatment Timing in California
A.8 Treatment at Varying Distances
@@
To determine shootings that are affected under the varying distance, we calculate a 5 and 25 mile radii, in euclidean distance, around each shooting’s location. We use the North American Datum of 1983 (NAD83) geodetic datum as our Coordinate Reference System (CRS). Figure A.8.1 shows these buffers, in blue and green, respectively. Figure A.8.1 zooms in on Southern California and shows shootings occurring between 2012 and 2022 only.
Figure A.8.1: Treatment Assignment Strategy for Varying Distances
@@
A.9 Assigning Treatment to a District
To estimate the effects of a mass shooting on the Gun Control Score of a district, we first confine our period of study to 2012-2022 to remove any concerns that we are not appropriately accounting for redistricting. Our estimation strategy is similar to those underlying our analyses of individual legislators, though we include chamber- and year-fixed effects and cluster standard errors by chamber and year. The full results can be seen in Table A.9.1.
Table A.9.1: ATT Estimates for Effect of Mass Shooting on District
@@ A.9 Assigning Treatment to a District
N 13,638 9,952 3,686
# of Treated Legislators 42 20 22
Chamber FE
Year FE
Pre-Treatment Mean 0.239 0.448 -0.233
  • Note: Robust standard errors are clustered at the chamber and year level. Gun Control Score is measured from -1 to 1, from least to most supportive of restrictive firearms legislation. Table reports the estimated effects of mass shootings occurring within a district on the Gun Control Score of the district. Individual legislators who switch parties throughout the duration of study are removed. Pre-Treatment Mean is the average gun control score of control and treated legislators in the pre-treatment period.
A.10 Cosponsorship Network Analysis and Gun Policy Activity
A.10.1 Cosponsorship Network
Figure A.10.1: Cosponsorship Network Among California Legislators on Firearm-related Legislation
Table A.10.1: Legislators with Highest Centrality in Each Chamber and Session
@@ A.10.1 Cosponsorship Network
A.10.2 Alternative Behavior
Table A.10.2: ATT Estimates for Effect of Mass Shooting on Legislators’ Activity in Gun Policy Domain
@@ A.10.2 Alternative Behavior
Legislator FE
Chamber FE
Year FE
Pre-Treatment Mean 0.016 0.018 0.010 0.799 0.964 0.397
N 7,924 5,898 2,026 7,924 5,898 2,026
  • Note: Standard errors are clustered at the legislator-, chamber-, and year level in parentheses. Table reports the estimated effects of mass shootings occurring within a legislator’s district on a legislator’s activity in the gun policy domain for periods $k=-2$ to 2. Activity is a measure of the number of bills sponsored or cosponsored by a legislator related to gun policy out of all bills related to gun policy in a given year. Outdegree is a measure of legislator’s level of activity in cosponsoring firearms-related legislation. Pre-Treatment Mean is the average gun control score of control and treated legislators in the pre-treatment period.